draw the graph of the equation x^2+y^2+4x+6y-3=0 and x^2+y^2+12x+14y+60=0. then find the equation of the radical axis and draw the axis.

Grade:

1 Answers

Vikas TU
11 years ago

since the eqn of the radical axis is S1 - S2 = 0

SUB tracting the eqn.s of the circles =

x^2+y^2+4x+6y-3 - (x^2+y^2+12x+14y+60) 

-8x -8y -63 = 0

we get, 8x + 8y + 63 = 0